### v1

# General properties of the electric Penrose process

Abstract
We consider the Penrose process with the charged particles in the Reissner-Nordström background. Let parent particle 0 decay to particles 1 and 2. With the assumption that all three particles move in the same plane, the exact formulas for characteristics of particles 1 and 2 in terms of those of particle 0 are derived. We concentrate on scenarios in which particle 1 and 2 are ejected along the trajectory of particle 0. It is shown that such scenarios correspond to the extrema of energies $E_{1}$ or $E_{2}$ of daughter particles with respect to the angular momentum $L_{1}$ or $L_{2}$ of one of them. We derive bounds on the values of angular momenta $L_{1}$ and $L_{2}$. We give classification of these scenarios and discuss their properties including decay in the near-horizon region. The results are reformulated in terms of velocities of daughter particles in the center of mass frame. The approach is applicable also to collisional Penrose process, in which a combination of particles 1 and 2 is considered as one effective particle. If the mass of particle 0 $m_{0}\rightarrow\infty$, the situation corresponds to the Bañados-Silk-West effect, the results agree with the ones known in literature before. In addition, we consider special cases when decay occurs in the turning point for one or all three particles.
energy extraction, rotating black hole
pacs:
04.70.Bw, 97.60.Lf
I Introduction
The Penrose process (hereafter PP) means that a parent particle 0 decays to two daughter ones 1 and 2 in such a way that one of them (say, 1) has a negative energy whereas particle 2 returns to infinity with an energy bigger than an initial energy $E_{0}$ of particle 0. This process becomes possible if in a space-time there exists the region (called ergoregion or ergosphere) inside which the component of the metric $g_{tt}>0$, if signature of a metric is chosen to be (-,+,+,+) pen , pen2 . Meanwhile, a counterpart of the original PP in the Reissner-Nordström (RN) background was found due to the properties of particle dynamics even in spite of the absence of such an ergoregion ruf , den . Instead, in this metric negative energies are possible in some region (called generalized ergoregion) whose border depends on the electric charge, angular momentum and mass of a particle.
At present, we see a new wave of interest to the PP in new contexts. This includes the collisional version of it schrev , confined one conf , myconf , the relation between the so-called Bañados-Silk-West (BSW) effect ban and the PP, discussion of the possibility of getting unbounded  energies called super-Penrose (SPP) effect (see, e.g. is and references therein), etc. As a rule, the investigations of the BSW effect were based on the careful analysis of the vicinity of the horizon. This implies that the value of the lapse function $N\ll 1$. Meanwhile, there is another approach that is based on exact formulas describing decay. For neutral particles this was realized in centr . The meaningful difference between (i) collisions near a rotating black hole and (ii) a static charged one consists in that in case (i) there is an upper bound forbidding the SPP near black holes but there is no such a bound for (ii) rn , fh .
The emphasis in ruf , den and the main part of subsequent works of this trend (see dad3 for review) was made on the properties of negative energy states for a given particle in the presence of the electromagnetic field for a concrete metric rufkerr . Meanwhile, we are going to focus on another aspect connected with the relation between an initial state and its products after decay. We elucidate how characteristics of particle 0 and those of particles 1 and 2 are related with each other, which scenarios are possible in principle and what are their properties. Although for a realistic process in astrophysics it is necessary to take into account both rotation and electromagnetic field that includes not only the electric but also the magnetic one dad1 , dad2 , luca , fifty (see also a useful list of references on the subject in win ), it looks reasonable at the first stage to consider a simplified problem and separate two different factors - rotation and electric charge. In the previous work eq we gave full classification of possible scenarios for the particle decay and PP with neutral particles in the rotating  background, provided all processes occur within the equatorial plane. In the present work, we consider a similar problem for electrically charged particles in the static black hole metric with the charge, i.e. the Reissner-Nodström (RN) one. It is worth noting that a similar problem was considered recently in ju but for weakly charged black holes only and one scenario whereas in our approach we take into account this charge exactly and give classification of all possible scenarios. What is more important, our results are qualitatively different (see discussion below).
We use system of units in which fundamental constants $G=c=1$.
The paper is organized as follows. In Sec. II we list equations of particle motion in the RN background. In Sec. III we give the exact formulas that related characteristics of particle 0 and particles 1, 2. In Sec. IV we suggest classification of scenarios when particle 0 is ingoing. In Sec. V we make emphasis on a type of decay when particles 1 and 2 are ejected along the trajectory of particle 0. In Sec. VI we consider a special case than an escaping particle is massless. In Sec. VII we consider some properties of scenario I including bounds on the energy and angular momentum of an escaping particle and decay in the near-horizon region. In Sec. VIII we discuss the situation when the point of decay coincides with the turning point of radial motion for particle 0, particle 2 or with a common turning point of all three particles. In Sec. IX we derive the conditions necessary for the PP process in the scenarios under discussion. In Sec. X we reformulate our result kinematically, using the velocities of particles and their Lorentz gamma-factors. In Sec. XI we consider the scenario in which particle 0 is outgoing. In Sec. XII we display the relation between our approach and that used in literature for the description of the BSW effect before. In Sec. XIII we give the summary of main results.
II Charged particles in the metric of Reissner-Nordström black hole
We consider the metric that has the form
$ds^{2}=-dt^{2}f+\frac{dr^{2}}{f}+r^{2}d\omega^{2},$ (1)
with $d\omega^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}.$ We will mainly deal with the RN metric, then $f\equiv N^{2}=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}=\frac{(r-r_{+})(r-r_{-})}{r^{2}}$. Here, $M$ is a black hole mass, $Q$ is its electric charge (for definiteness $Q>0$), $r_{+}=M+\sqrt{M^{2}-Q^{2}}$ being the event horizon radius, $r_{-}$ the inner horizon radius. Meanwhile, basic results are extendable to an arbitrary $f$ outside the event horizon.
Let us consider motion of a charged particle in this space-time without additional forces. Then, there are two integrals of motion. As the metric does not depend on time and angle $\phi$, the energy $E$ and angular momentum $L$ are conserved. Motion occurs within a plane which we choose to be $\theta=\frac{\pi}{2}$. Then, equations of motion read (dot denotes derivative with respect to the proper time $\tau$)
$\dot{t}=\frac{X}{mf}\text{,}$ (2)
$X=E-q\varphi\text{,}$ (3)
$\dot{\phi}=\frac{L}{mr^{2}}\text{,}$ (4)
$\dot{r}=\sigma P\text{,}$ (5)
$P=\sqrt{X^{2}-N^{2}\tilde{m}^{2}}\text{,}$ (6)
$\tilde{m}^{2}=m^{2}+\frac{L^{2}}{r^{2}}.$ (7)
Here, $q$ is a particle’s charge, $\varphi=\frac{Q}{r}$ is the Coulomb potential of the RN metric , $\sigma=\pm 1$. The forward-in-time condition $\dot{t}>0$ requires
$X\geq 0\text{,}$ (8)
outside the horizon $X>0$.
From $P^{2}\geq 0$ we have a more tight condition than (8):
$X\geq\tilde{m}N.$ (9)
III Decay to two particles: general scenario
Let in some point $r=r_{d}$ particle 0 decay to particles 1 and 2. For simplicity, we assume that all three particles move within the same plane. In the point of decay the conservation laws give us
$E_{0}=E_{1}+E_{2},$ (10)
$L_{0}=L_{1}+L_{2}\text{,}$ (11)
$q_{0}=q_{1}+q_{2}\text{.}$ (12)
It follows from (10) - (12) that
$X_{0}=X_{1}+X_{2}.$ (13)
The conservation of the radial component of momentum reads
$\sigma_{0}P_{0}=\sigma_{1}P_{1}+\sigma_{2}P_{2}\text{.}$ (14)
Let particle 0 move with decreasing $r$, so $\sigma_{0}=-1$. It is clear that combination $\sigma_{1}=\sigma_{2}=+1$ is impossible since this would contradict eq. (14). For definiteness, we assume that particle 1 moves after decay with $\sigma_{1}=-1$. Moreover, we imply that in the frame comoving with particle 0 (which is the center of mass frame of particles 1 and 2 in the point of decay), particle 1 moves in the inward direction as well. Meanwhile, $\sigma_{2}$ can have any sign. In the center of mass frame particle 2 moves in the outward direction with some velocity $v_{2}$. However, if we pass to the frame comiving with particle 0 that moves in the inward direction with the velocity $V_{2}$ in the static frame, both cases are possible: $\sigma_{2}=-1$ or $\sigma_{2}=+1$ depending on which value of the velocity is bigger (below we will discuss this issue in more detail).
Then we have
$P_{0}=P_{1}-\sigma_{2}P_{2}\text{.}$ (15)
Solving equations of motion, we obtain
$X_{1}=\frac{X_{0}}{2\tilde{m}_{0}^{2}}\tilde{b}_{1}-\delta\frac{P_{0}\sqrt{\tilde{d}}}{2\tilde{m}_{0}^{2}}\text{,}$ (16)
$X_{2}=\frac{1}{2\tilde{m}_{0}^{2}}\left(X_{0}\tilde{b}_{2}+P_{0}\delta\sqrt{\tilde{d}}\right),$ (17)
$P_{1}=\left|\frac{P_{0}\tilde{b}_{1}-\delta X_{0}\sqrt{\tilde{d}}}{2\tilde{m}_{0}^{2}}\right|\text{.}$ (18)
$P_{2}=\left|\frac{P_{0}\tilde{b}_{2}+\delta X_{0}\sqrt{\tilde{d}}}{2\tilde{m}_{0}^{2}}\right|.$ (19)
Here, $\delta=\pm 1$,
$\tilde{b}_{1,2}=\tilde{m}_{0}^{2}+\tilde{m}_{1,2}^{2}-\tilde{m}_{2,1}^{2}\text{,}$ (20)
$\tilde{d}=\tilde{b}_{1}^{2}-4\tilde{m}_{0}^{2}\tilde{m}_{1}^{2}=\tilde{b}_{2}^{2}-4\tilde{m}_{0}^{2}\tilde{m}_{2}^{2}\text{.}$ (21)
For what follows, it is also convenient to introduce also corresponding quantities without tilde:
$b_{1,2}=m_{0}^{2}+(m_{1,2}^{2}-m_{2,1}^{2}),$ (22)
$d=b_{1}^{2}-4m_{0}^{2}m_{1}^{2}=b_{2}^{2}-4m_{0}^{2}m_{2}^{2}.$ (23)
Then, it can be obtained from (11) that
$\tilde{b}_{1,2}=b_{1,2}+\frac{2L_{0}L_{1,2}}{g_{\phi}},$ (24)
$\tilde{d}=d+4b_{1}\frac{L_{0}L_{1}}{r^{2}}-\frac{4L_{1}^{2}}{r^{2}}m_{0}^{2}-4\frac{L_{0}^{2}}{r^{2}}m_{1}^{2},$ (25)
$\tilde{d}=d+4b_{2}\frac{L_{0}L_{2}}{r^{2}}-\frac{4L_{2}^{2}}{r^{2}}m_{0}^{2}-4\frac{L_{0}^{2}}{r^{2}}m_{2}^{2}.$ (26)
From the condition $\tilde{d}\geq 0$ one can deduce that
$\tilde{m}_{0}\geq\tilde{m}_{1}+\tilde{m}_{2}.$ (27)
Eqs. (16) - (19) are the counterpart of equations (19), (20), (26), (27) in centr obtained there for rotating metrics. The essential difference consists in the expression for $X$. Now it is given by (3) and does not depend on $L$. Meanwhile, in centr $X=E-\omega L$, the $\omega$ is the metric coefficient responsible for rotation. It is the fact that $L$ does not enter the expression for $X$ that allows some scenarios which were forbidden in the rotating case (see below).
IV Classification of scenarios in general
Each scenario of decay can be characterized by the set $(\sigma_{2}$, $h_{2}$, $h_{1}$, $\delta)$, where
$h_{1,2}=sign(\tilde{b}_{1,2}N-2\tilde{m}_{1,2}X_{0}).$ (28)
The quantities $h_{1}$ and $h_{2}$ arise in a natural way, when we sort out combinations with different signs inside the absolute values in (18) and (19) and determine which term dominates. As a result, one obtains the following types of scenarios:
I $(\sigma_{2}=+1$, $h_{2}\geq 0$, $\delta=-1)$
$P_{1}-P_{2}=P_{0}\text{,}$ (29)
$P_{2}=\frac{(X_{0}\sqrt{\tilde{d}}-P_{0}\tilde{b}_{2})}{2\tilde{m}_{0}^{2}}\text{,}$ (30)
$P_{1}=\frac{P_{0}\tilde{b}_{1}+X_{0}\sqrt{\tilde{d}}}{2\tilde{m}_{0}^{2}},$ (31)
$X_{1}=\frac{X_{0}}{2\tilde{m}_{0}^{2}}\tilde{b}_{1}+\frac{P_{0}\sqrt{\tilde{d}}}{2\tilde{m}_{0}^{2}}\text{,}$ (32)
$X_{2}=\frac{1}{2\tilde{m}_{0}^{2}}\left(X_{0}\tilde{b}_{2}-P_{0}\sqrt{\tilde{d}}\right).$ (33)
II $(\sigma_{2}=-1$, $h_{2}<0$, $\delta=-1)$
$P_{1}+P_{2}=P_{0},$ (34)
$P_{2}=\frac{(P_{0}\tilde{b}_{2}-X_{0}\sqrt{\tilde{d}})}{2\tilde{m}_{0}^{2}}\text{,}$ (35)
$P_{1}=\frac{P_{0}\tilde{b}_{1}+X_{0}\sqrt{\tilde{d}}}{2\tilde{m}_{0}^{2}}\text{,}$ (36)
$X_{1}=\frac{X_{0}}{2\tilde{m}_{0}^{2}}\tilde{b}_{1}+\frac{P_{0}\sqrt{\tilde{d}}}{2\tilde{m}_{0}^{2}}\text{,}$ (37)
$X_{2}=\frac{1}{2\tilde{m}_{0}^{2}}\left(X_{0}\tilde{b}_{2}-P_{0}\delta\sqrt{\tilde{d}}\right)\text{.}$ (38)
III $(\sigma_{2}=-1$, $h_{1}\leq 0$, $\delta=+1)$
$P_{2}=\frac{X_{0}\sqrt{\tilde{d}}+P_{0}\tilde{b}_{2}}{2\tilde{m}_{0}^{2}}\text{,}$ (39)
$P_{1}=\frac{P_{0}\tilde{b}_{1}-X_{0}\sqrt{\tilde{d}}}{2\tilde{m}_{0}^{2}}\text{,}$ (40)
$X_{1}=\frac{X_{0}}{2\tilde{m}_{0}^{2}}\tilde{b}_{1}-\frac{P_{0}\sqrt{\tilde{d}}}{2\tilde{m}_{0}^{2}}\text{,}$ (41)
$X_{2}=\frac{1}{2\tilde{m}_{0}^{2}}\left(X_{0}\tilde{b}_{2}+P_{0}\sqrt{\tilde{d}}\right).$ (42)
Scenarios II and III can be obtained from each other by interchange of particles 1 and 2. In this sense, they are equivalent, so for definiteness we will consider scenario II.
The condition $\tilde{d}\geq 0$ requires that, according to (25), (26)
$L_{1,2}^{(-)}\leq L_{1,2}\leq L_{1,2}^{(+)}\text{,}$ (43)
where
$L_{1,2}^{(\pm)}=\frac{b_{1,2}L_{0}}{2m_{0}^{2}}\pm\frac{\sqrt{d}r_{d}}{2m_{0}^{2}}\tilde{m}_{0}\text{.}$ (44)
V Ejection along trajectory: maximization/minimization of energy and classification of scenarios
Now, we consider a special type of decay in which products of decay are ejected along the trajectory of the original particle.
In doing so,
$\frac{P_{1}}{P_{0}}=\frac{L_{1}}{L_{0}}\text{,}$ (45)
whence
$(P_{0}\tilde{b}_{1}-\delta X_{0}\sqrt{\tilde{d}})L_{0}-2\tilde{m}_{0}^{2}L_{1}P_{0}=0\text{.}$ (46)
For definiteness, we assumed that signs of angular momenta of particles 0 and 1 coincide.
The interest to such a type of trajectory is motivated by the fact that it is the most efficient process (with other parameters of scenario fixed). The particle that moves in the same direction as particle 0 acquires the maximum possible energy, the other one that moves in the opposite direction has the minimum one. To show this, one can consider, say, energy $E_{1}$ as a function of $L_{1}$:
$E_{1}=\frac{X_{0}}{2\tilde{m}_{0}^{2}}\tilde{b}_{1}-\delta\frac{P_{0}\sqrt{\tilde{d}}}{2\tilde{m}_{0}^{2}}+q_{1}\varphi\text{.}$ (47)
In the case of extremum, we have
$\frac{\partial E_{1}}{\partial L_{1}}=0\text{.}$ (48)
After taking the square and performing a number of algebraic manipulations, we obtain just expressions for $X_{1}$ and $X_{2}$ that can be obtained from (45). We list them below explicitly.
There are two different possible scenarios here. We consider them separately.
V.1 Scenario I, $\sigma_{2}=+1$, $\delta=-1$, $h_{2}=+1$
Remarkably, it turns out that after somewhat long algebraic manipulation, one can get rid off tilted quantities and express the results in a rather simple form in terms of quantities without tilde:
$X_{2}=X_{0}\frac{b_{2}}{2m_{0}^{2}}-\frac{\sqrt{d}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}}{2m_{0}^{2}},$ (49)
$X_{1}=X_{0}\frac{b_{1}}{2m_{0}^{2}}+\frac{\sqrt{d}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}}{2m_{0}^{2}}.$ (50)
$P_{2}=\frac{P_{0}}{2m_{0}^{2}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}}(X_{0}\sqrt{d}-b_{2}\sqrt{X_{0}^{2}\mathcal{-}N^{2}m_{0}^{2}})\text{,}$ (51)
$P_{1}=\frac{P_{0}}{2m_{0}^{2}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}}(b_{1}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}+X_{0}\sqrt{d}).$ (52)
In general in eqs. (16) - (19) one of angular momenta (say, $L_{1}$) was a free parameter. However, now this is not so since there is an additional constraint (45) that leads to
$L_{2}=\frac{L_{0}}{2m_{0}^{2}}(b_{2}-\frac{\sqrt{d}}{\sqrt{X_{0}^{2}-N^{2}m_{0}^{2}}}X_{0})\text{,}$ (53)
$L_{1}=\frac{L_{0}}{2m_{0}^{2}}(b_{1}+\frac{\sqrt{d}X_{0}}{\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}})\text{.}$ (54)
One can check that eqs.  (6) and (45) do hold for each particle 1 and 2.
The requirement $d\geq 0$ entails the inequality
$m_{0}\geq m_{1}+m_{2},$ (55)
similar to (27).
It follows from $P_{2}\geq 0$ that
$b_{2}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}\leq X_{0}\sqrt{d}\,,$ (56)
whence
$X_{0}\leq\frac{Nb_{2}}{2m_{2}}.$ (57)
As the expression inside the square root should be non-negative, $X_{0}\geq m_{0}N$. It holds automatically, if (9) does so. Combining the above inequalities and (9), we have
$\tilde{m}_{0}N\leq X_{0}\leq\frac{Nb_{2}}{2m_{2}}.$ (58)
For $L_{0}=0$ (so $\tilde{m}_{0}=m_{0}$) this is consistent with (57) automatically since
$b_{2}\geq 2m_{0}m_{2}$ (59)
due to (55).
V.2 Scenario II
$\sigma_{2}=-1\,,$ $\delta=-1$
$P_{1}+P_{2}=P_{0\text{,}}$ (60)
$L_{1}=\frac{L_{0}}{2m_{0}^{2}}(b_{1}+\frac{\sqrt{d}X_{0}}{\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}})\text{,}$ (61)
$L_{2}=\frac{L_{0}}{2m_{0}^{2}}(b_{2}-\frac{\sqrt{d}X_{0}}{\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}}),$ (62)
$X_{1}=\frac{X_{0}b_{1}}{2m_{0}^{2}}+\frac{\sqrt{d}}{2m_{0}^{2}}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}},$ (63)
$X_{2}=\frac{b_{2}X_{0}}{2m_{0}^{2}}-\frac{\sqrt{d}}{2m_{0}^{2}}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}.$ (64)
$P_{1}=\frac{P_{0}}{2m_{0}^{2}\sqrt{X_{0}^{2}-N^{2}m_{0}^{2}}}(b_{1}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}+X_{0}\sqrt{d})\text{,}$ (65)
$P_{2}=\frac{P_{0}}{2m_{0}^{2}\sqrt{X_{0}^{2}-N^{2}m_{0}^{2}}}(b_{2}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}-X_{0}\sqrt{d}).$ (66)
It follows from $P_{2}\geq 0$ that
$X_{0}\geq\frac{b_{2}N}{2m_{2}}$ (67)
and
$Nm_{2}\leq X_{2}\leq\frac{b_{2}}{2m_{0}^{2}}X_{0}\text{,}$ (68)
where we took into account (67) and (22), (23).
It is worth noting that in scenario I $L_{2}$ has the sign opposite to that of $L_{0}$, in scenario II they coincide.
If $Q\rightarrow 0$, we arrive at the Schwarzschild metric. Then, our formulas agree with ob . In this case, there is no the PP and the energy of an escaping particle $E_{2}<E_{0}$ but $E_{2}\neq 0$.
VI Massless case
A special case arises when particle 2 is massless since condition (67) is no longer valid if $m_{2}=0$. In turn, this means that scenario II fails. This can be explained as follows. In general, scenario II is realized when particle 2 is ejected in the outward direction in the frame comoving with particle 0 (that is also the center of mass frame for particles 1 and 2). However, when we pass to the stationary frame, it is drifted in the inward direction due to motion of particle 0 (see Sec. X below for discussion of local Lorentz transformations). However, if particle 2 is massless, it moves with a speed of light and inward motion of particle 0 cannot overcome outward motion of particle 2. As a result, particle 2 moves outwardly that corresponds to scenario I, not II.
Therefore, we must use eqs. (49), (51), in which we put $m_{2}=0$ that gives us $b_{2}=d=m_{0}^{2}-m_{1}^{2}$. Thus we have
$X_{2}=\frac{b_{2}}{2m_{0}^{2}}(X_{0}-\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}),$ (69)
$P_{2}=\frac{P_{0}b_{2}}{2m_{0}^{2}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}}(X_{0}-\sqrt{X_{0}^{2}\mathcal{-}N^{2}m_{0}^{2}})\text{.}$ (70)
From another hand, if we put $m_{1}=0$, this does not give rise to any difficulties since particle 1 moves inwardly in both frames, so scenario II is valid.
In the horizon limit $N\rightarrow 0$,
$X_{2}\approx\frac{b_{2}N^{2}}{4X_{0}}\text{, }E\approx\frac{q_{2}Q}{r_{+}}+\frac{b_{2}N^{2}}{4X_{0}}\text{,}$ (71)
$X_{1}\approx X_{0}-\frac{b_{2}N^{2}}{4X_{0}}\text{, }E_{1}\approx E_{0}-\frac{q_{2}Q}{r_{+}}-\frac{b_{2}N^{2}}{4X_{0}}\text{.}$ (72)
VII Properties of scenario I
We are interested mainly in scenario I since it is this scenario in which particle 2 can escape to infinity. First of all, we would like to stress the difference between scenario I and its counterpart for processes with neutral particles in the background of rotating black holes. In the latter case, decay of type I is impossible in the ergoregion at all eq . Meanwhile, now it is allowed inside a generalized ergoregion. Below, we discuss some details concerning just scenario I.
VII.1 Upper bound on the angular momentum
For scenario I, there are restrictions (58). Both inequalities here are consistent with each other, provided
$\tilde{m}_{0}\leq\frac{b_{2}}{2m_{2}}\text{.}$ (73)
Taking the square, we obtain the restriction on a value of $L_{0}$:
$L_{0}^{2}\leq L_{m}^{2}=\frac{dr_{d}^{2}}{4m_{2}^{2}}\text{.}$ (74)
Let $d\rightarrow 0$. This entails
$m_{0}=m_{1}+m_{2}.$ (75)
Then, scenario I is possible if $L_{0}\rightarrow 0$ only. According to (53), (54) this entails that $L_{1,2}\rightarrow 0$ as well, so all three particles move radially. This scenario is realized in the confined Penrose process  conf , myconf . There is no similar restriction on $L_{0}$ in scenario II.
VII.2 Upper bound on the energy of escaping particle
It follows from (57) that scenario I can be realized for $\left(X_{0}\right)_{d}\leq\frac{N_{d}b_{2}}{2m_{2}}$ only. Then, from (49) and (22), (23) we have
$Nm_{2}\leq X_{2}\leq\frac{b_{2}^{2}N}{4m_{0}^{2}m_{2}}.$ (76)
Thus for the escaping particle we have the upper bounds
$q_{2}\frac{Q}{r_{d}}+Nm_{2}\leq E_{2}\leq q_{2}\frac{Q}{r_{d}}+\frac{b_{2}^{2}N}{4m_{0}^{2}m_{2}}\text{.}$ (77)
VII.3 Near-horizon limit
Assuming that $m_{2}\neq 0$, let us consider
$\left(X_{0}\right)_{d}=a\frac{N_{d}b_{2}}{2m_{2}}$ (78)
where according to (58) the coefficient $a$ (pure number) obeys inequalities
$a_{m}\leq a\leq 1\text{, }a_{m}=2\frac{m_{2}\tilde{m}_{0}}{b_{2}}$ (79)
in combinations with inequalities (73), (74).
If $N_{d}\rightarrow 0$ while keeping $a$ constant, we obtain an approximate equality:
$E_{2}\approx q_{2}\frac{Q}{r_{+}}+cm_{2}N_{d}\text{,}$ (80)
where
$c=\frac{b_{2}^{2}a-\sqrt{d}\sqrt{a^{2}b_{2}^{2}-4m_{0}^{2}m_{2}^{2}}}{4m_{2}^{2}m_{0}^{2}}\text{.}$ (81)
One can check that $\frac{\partial c}{\partial a}\leq 0$. Therefore,
$1\leq c\leq c(a_{m})=c_{m}\text{,}$ (82)
where
$c_{m}=\frac{\tilde{m}_{0}(r_{+})b_{2}}{2m_{2}m_{0}^{2}}-\frac{\left|L_{0}\right|\sqrt{d}}{2m_{2}m_{0}^{2}r_{+}}.$ (83)
In particular, if $L_{0}=0$,
$c_{m}=\frac{b_{2}}{2m_{2}m_{0}}\geq 1$ (84)
that agrees with (77). If $L_{0}=L_{m}$, we obtain from (74)
$c_{m}=1.$ (85)
If $a=1$, it follows that $c=1$ and we return to (77) with equality instead of inequality,   so $E_{2}\approx q_{2}\frac{Q}{r_{+}}+m_{2}N_{d}.$
Thus the correction to the main term in $X_{2}$ has the order $N$. We arrive at an important conclusion. If decay occurs near the horizon, the particle that moves to infinity should be near-critical with $X_{0}=O(N_{d})$. Usual particles cannot escape from the horizon at all.
In the massless case $m_{2}=0$ the correction has the order $N^{2}$ according to (71).
For particle 1 we have in the near-horizon limit
$E_{1}\approx q_{1}\frac{Q}{r_{+}}+N_{d}(\frac{ab_{2}}{2m_{2}}-c)\text{.}$ (86)
The Penrose process with charged particles was also considered in ju where an additional assumption that the black hole charge $Q$ is small was made. Our results do not agree. The authors of ju , according to their eqs. (34), (35), obtained that the outgoing particle after decay near the horizon has a finite nonzero energy. Meanwhile, it follows from our formulas that in this case the energy of an escaping particle (if we neglect $Q$) tends to zero. If we take the charge $Q$ into account, in a similar way the difference $X_{2}=E_{2}-q_{2}\frac{Q}{r_{+}}$ tends to zero. Actually, the subtlety in the issue under discussion consists in the necessity to take into account correlation between dynamics and kinematics. This means that in scenario I (where particle 2 escapes) the sign before the radical in (49) comes with minus, not plus. Instead, one can consider particle 1 and finite $X_{1}$ with the sign plus but this particle falls in a black hole and does not escape.
One more attempt of considering the PP for charged particles was made in sh with a magnetic field $B$ taken into account. In our view, the results for efficiency $\eta=\frac{E_{2}}{E_{0}}$ m (in our notations) described by Eqs. (31) and (32) of Ref. sh are incorrect. In the neutral case $q=0$, $Q=0$ and $B=0$ they give $\eta=0$ instead of the Schwarzschild value $\eta_{Sch}$. Therefore, for the RN metric, it also does not reproduce the formulas like (106) of the present paper.
VIII Special type of scenario: decay in turning point
There are special cases to be considered separately. They arise when decay occurs in the turning point. (Hereafter, by turning point we imply for brevity a turning point for radial motion, a particle can have in general an angular moment and nonzero angular velocity.) It follows from the conservation law (14) that the point of decay cannot coincide with the turning point for two particles precisely. Either it is the turning point for (i) only one particle or (ii) for all three at once. It follows from (14) or (52), (65) that in scenario I, case (i) is possible in the case of the turning point for particle 2, not for particle 1. Now, we will discuss (i) and (ii) case by case.
VIII.1 Case (i), turning point for particle 2 (TP2)
Let particle 2 arise in its own turning point, so $P_{2}=0$ but $P_{0}\neq 0$. We call it scenario TP2. It follows from (16) - (19) that
$X_{2}=\frac{2\tilde{m}_{2}^{2}X_{0}}{\tilde{b}_{2}}\text{,}$ (87)
$X_{1}=\frac{X_{0}}{\tilde{b}_{2}}(\tilde{m}_{0}^{2}-\tilde{m}_{1}^{2}-\tilde{m}_{2}^{2})\text{,}$ (88)
$X_{0}=\frac{\tilde{b}_{2}N}{2\tilde{m}_{2}}\text{.}$ (89)
For ejection along the trajectory, the case under discussion is realized in scenario II and we obtain from (51), (66) with $P_{2}=0$, $P_{0}\neq 0$ that
$X_{0}=\frac{b_{2}N}{2\tilde{m}_{2}}\text{.}$ (90)
$X_{2}=m_{2}N=\frac{2m_{2}^{2}}{b_{2}}X_{0}.$ (91)
Then, (13) gives us
$X_{1}=\frac{X_{0}}{b_{2}}(m_{0}^{2}-m_{1}^{2}-m_{2}^{2})\text{.}$ (92)
From (53), (62) we have
$L_{2}=0\text{, }L_{1}=L_{0}\text{.}$ (93)
VIII.2 Case (i), turning point for particle 0 (TP0)
In general scenario we take the limit $P_{0}\rightarrow 0$. This means that decay occurs in the turning point of particle 0. For shortness, we call it TP0. Particles 1 and 2 may have nonzero $P_{1}$ and $P_{2}$, so the point $r_{d}$ is not the turning point for them. It is clear from the conservation law (14) that scenario TP0 can be realized in scenario I but not in II. We obtain from (16) - (19) that
$P_{1}=P_{2}=\frac{N\sqrt{\tilde{d}}}{2\tilde{m}_{0}},$ (94)
$X_{1}=\frac{N}{2\tilde{m}_{0}}\tilde{b}_{1},$ (95)
$X_{2}=\frac{N}{2\tilde{m}_{0}}\tilde{b}_{2}.$ (96)
Three quantities $E_{0}$, $L_{0}$, $N$ are related by one equation
$X_{0}=\tilde{m}_{0}N\text{.}$ (97)
It follows from (97) immediately that
$L_{0}^{2}=r_{d}^{2}\frac{X_{0}^{2}-m_{0}^{2}N^{2}}{N^{2}}\text{.}$ (98)
In general, $L_{1}$ (or $L_{2}=L_{0}-L_{1})$ is a free parameter. It is only restricted by the condition (43).
Let particles 1 and 2 be ejected in point $r_{d}$ along the trajectory of particle 0. Now, it follows from (45) that either $L_{0}=0$ or $P_{1}=0=P_{2}$. In the first case, $\tilde{m}_{0}=m_{0}$, so
$X_{0}=m_{0}N.\text{ }$ (99)
Then, it follows from (49) - (54) that
$X_{1}=X_{0}\frac{b_{1}}{2m_{0}^{2}}=\frac{b_{1}}{2m_{0}}N\text{,}$ (100)
$X_{2}=X_{0}\frac{b_{2}}{2m_{0}^{2}}=\frac{b_{2}}{2m_{0}}N\text{.}$ (101)
In the second case the point $r_{d}$ is the turning point for all three particles and we turn to this possibility in the next subsection.
VIII.3 Case (ii), turning point for all three particles (TP3)
Let us call this scenario TP3. Eq. It follows from (6) that
$X_{0}=\tilde{m}_{0}N\text{,}$ (102)
$X_{1}=\tilde{m}_{1}N\text{, }$ (103)
$X_{2}=\tilde{m}_{2}N\text{.}$ (104)
With (13) taken into account, this leads to
$\tilde{m}_{0}=\tilde{m}_{1}+\tilde{m}_{2}\text{.}$ (105)
Then, after some algebraic manipulations, we can again obtain the same expressions for $X_{1}$ and $X_{2}$ as in scenarios I and II, but with one important difference. In scenario I, the expression for $X_{2}$ and, correspondingly, energy $E_{2}$ for particle 2 that is enable to escape to infinity, contained sign ”minus” before the square root. This was due to the necessity to have non-negative factor in (51) inside parentheses. Meanwhile, now this is irrelevant since $P_{2}=0$ due to the factor $P_{0}=0$. As a result, escaping particle 2 can have not only sign ”minus” but also sign ”plus”, so we are free to take
$E_{2}=q_{2}\varphi+\frac{b_{2}}{2m_{0}^{2}}X_{0}+\frac{\sqrt{d}}{2m_{0}^{2}}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}\text{.}$ (106)
If so, for particle 1 we have
$E_{1}=q_{1}\varphi+\frac{b_{1}}{2m_{0}^{2}}X_{0}-\frac{\sqrt{d}}{2m_{0}^{2}}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}\text{.}$ (107)
Equivalently,
$E_{2}=q_{2}\varphi+\frac{b_{2}}{2m_{0}^{2}}\tilde{m}_{0}N+\frac{\sqrt{d}\left|L_{0}\right|}{2m_{0}^{2}r_{d}}\text{,}$ (108)
$E_{1}=q_{1}\varphi+\frac{b_{1}}{2m_{0}^{2}}\tilde{m}_{0}N-\frac{\sqrt{d}\left|L_{0}\right|}{2m_{0}^{2}r_{d}}\text{.}$ (109)
From (105) we have, taking the square and solving the quadratic equation that
$L_{2}=\frac{1}{2m_{0}^{2}}(L_{0}b_{2}+signL_{0}\sqrt{d}\tilde{m}_{0}r_{d})=L_{2}^{(+)},$ (110)
$L_{1}=\frac{1}{2m_{0}^{2}}(L_{0}b_{1}-signL_{0}\sqrt{d}\tilde{m}_{0}r_{d})=L_{1}^{(-)}\text{,}$ (111)
where $L^{(+)}$, $L^{(-)}$ are defined in (44). Distinction between particles 1 and 2 is almost conditional now but with the reservation that the sign in the second term in (108) or in (109) correlates with that in (110), (111).
In a sense, when from scenarios I and II we pass to scenario TP3, there is an exchange of branches ”plus” and ”minus” between both particles in the point where $P_{0}=0$ for non-radial motion. The aforementioned difference disappears if $L_{0}=0$. Then, it follows from (102) that $X_{0}=m_{0}N$. If also $d=0$, we return to the process described in conf , myconf . However, for nonzero $L_{0}$ the case under discussion is more general.
When both $P_{0}\rightarrow 0$ and $\tilde{d}\rightarrow 0$, all scenarios discussed in this section agree with each other.
IX Threshold for the Penrose process
Now, we are going to elucidate, when the Penrose process is possible. For rotating black holes, this requires the existence of the ergoregion where, by definition, $g_{00}>0$ pen . In the RN case, $g_{00}$ does not change the sign outside the horizon and there is no ergoregion in a usual sense but there exists its analogue - so-called generalized ergosphere ruf , den where negative energy are allowed. It is sensitive to the properties of particles. When $E_{1}<0\,$, the PP becomes possible and particle 2 with an excess of energy goes to infinity. Meanwhile, in scenario II both particles fall in a black hole. Therefore, we consider scenario I. The condition $E_{1}<0$ leads to
$X_{0}\frac{b_{1}}{2m_{0}^{2}}+q_{1}\varphi+\frac{\sqrt{d}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}}{2m_{0}^{2}}\leq 0.$ (112)
For $q_{1}\geq 0$ this is impossible. Let $q_{1}=-\left|q_{1}\right|<0$. Then this condition takes the form
$2\left|q_{1}\right|m_{0}^{2}\varphi\geq X_{0}b_{1}+\sqrt{d}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}.$ (113)
Moreover, the SPP is now possible as well. Indeed, for a fixed $q_{0}$, the energy $E_{2}=X_{2}+q_{2}\varphi$ formally grows unbounded when $q_{2}\rightarrow\infty$. The fact that the electromagnetic field can significantly enhance the efficiency of the Penrose process was pointed in dad1 , dad2 . Meanwhile, for sufficiently high $q_{2}$ this is quite generic feature that does not require the presence of a black hole. This can happen even in the flat space-time flat .
Let us consider an important particular case. If, in the framework of scenario I, decay occurs in the turning point of particle 2, we can substitute here (90) in the point of decay and obtain
$E_{2}=q_{2}\varphi+N\frac{(b_{2}^{2}-d)}{4m_{0}^{2}m_{2}}\text{,}$ (114)
$E_{1}=q_{1}\varphi+N\frac{(b_{1}b_{2}+d)}{4m_{0}^{2}m_{2}}=q_{1}\varphi+N\frac{m_{0}^{2}-m_{1}^{2}-m_{2}^{2}}{2m_{2}}.$ (115)
It follows from (115) and (22), (23) that in this case (113) can be rewritten in the form
$\left|q_{1}\right|\varphi>N\frac{m_{0}^{2}-m_{1}^{2}-m_{2}^{2}}{2m_{2}}\text{.}$ (116)
If all $q_{i}\rightarrow 0$, the results coincide with those obtained in the static limit of a rotating black hole - see eq. 111 in ob , wald , eq. 3.30 in pani and eq and . Then, the Penrose process is impossible as it should be. For $\left|q_{1}\right|\neq 0$, the PP is possible in the scenario under discussion, if the point of decay, according to (116), is located sufficiently close to the horizon.
In the particular case, when $d=0$, we have $m_{0}=m_{1}+m_{2}$ conf and the above condition turns into
$\left|q_{1}\right|\varphi>N_{d}m_{1}\text{,}$ (117)
typical of the PP for pure radial motion, if decay occurs in the turning point conf .
Eq. (117) can be interpreted as the statement that the electrostatic energy of particle in an external fled should be bigger than the red-shifted energy measured by a local observer. It is worth noting that in the case under discussion $r_{d}$ is the turning point for particle 2, but not for particles 0 and 1. In particular, eq. (75) is not satisfied in general.
It is instructive to note, for comparison, that in the case of rotating metrics extraction of energy in scenario I inside the ergoregion cannot be realized at all (see eq , Sec. VIA).
X Velocities and gamma factors for scenario I
The above results are given in terms of particles’ masses. Meanwhile, the approach can be reformulated kinematically in terms of velocities. For definiteness, let us consider scenario I, so $\sigma_{2}=+1$. Then, one can check using (125) and (49) the validity of the equation
$V_{2}=\frac{v_{2}-V_{0}}{1-v_{2}V_{0}}$ (118)
that is nothing but the Lorentz law of adding velocities. Now, the gamma factor of relative motion for particles 0 and 2 $\gamma_{02}=-u_{0\mu}u_{2}^{\mu}$, whence
$m_{2}m_{0}\gamma_{02}=\frac{X_{0}X_{2}+P_{0}P_{2}}{N^{2}}-\frac{L_{0}L_{2}}{r^{2}}\text{.}$ (119)
In (119) we used eqs. of motion (2) - (6) and took into account that both particles move in the opposite direction that gives us sign ”plus” in the numerator of eq. (119). Then, one obtains
$\gamma_{02}=\frac{b_{2}}{2m_{0}m_{2}}\text{, }v_{2}=\frac{\sqrt{d}}{b_{2}}\text{.}$ (120)
Thus we can rewrite the formula for $X_{2}$ in the form
$X_{2}=\frac{m_{2}}{m_{0}}\gamma_{02}(X_{0}-v_{2}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}).$ (121)
In a similar manner, we find
$m_{2}m_{0}\gamma_{01}=\frac{X_{0}X_{1}-P_{0}P_{1}}{N^{2}}-\frac{L_{0}L_{1}}{r^{2}},$ (122)
$\gamma_{01}=\frac{b_{1}}{2m_{0}m_{1}}\text{, }v_{1}=\frac{\sqrt{d}}{b_{1}}\text{.}$ (123)
$X_{1}=\frac{m_{1}}{m_{0}}\gamma_{01}(X_{0}+v_{1}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}).$ (124)
Here, $\gamma_{0i}$ ($i=1,2$) has the meaning of the standard Lorentz factor of relative motion. For an individual particle we have
$X_{i}=m\gamma_{i}N\text{, }\gamma_{i}=\frac{1}{\sqrt{1-V_{i}^{2}}}\text{, }i=0,1,2\text{,}$ (125)
so
$V_{i}=\sqrt{1-\left(\frac{m_{i}N}{X_{i}}\right)^{2}\text{,}}$ (126)
$V_{i}$ is the velocity is measured in the stationary frame.
The expressions for the energy can be rewritten as
$E_{2}=q_{2}\varphi+\gamma_{02}\frac{m_{2}}{m_{0}}(X_{0}-v_{2}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}})\text{.}$ (127)
$E_{1}=q_{1}\varphi+\gamma_{01}\frac{m_{1}}{m_{0}}(X_{0}+v_{1}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}})\text{.}$ (128)
For the existence of the PP, the key restriction on the electric charge (113) is required. It can be written in terms of velocities
$\frac{m_{1}}{m_{0}}(X_{0}+v_{1}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}})<\left|q_{1}\right|\varphi\sqrt{1-v_{1}^{2}}\text{,}$ (129)
whence
$1+v_{1}V_{0}<\rho\sqrt{1-v_{1}^{2}}\text{, }\rho=\frac{m_{0}}{X_{0}m_{1}}\left|q_{1}\right|\varphi\text{. }$ (130)
Using (125), we can also write $\rho=\frac{\left|q_{1}\right|\varphi}{m_{1}N\gamma_{0}}$. Taking the square of (130), we obtain
$v_{1}^{2}+\frac{2v_{1}V_{0}}{V_{0}^{2}+\rho^{2}}+\frac{1-\rho^{2}}{V_{0}^{2}+\rho^{2}<0}=(v_{1}-v_{+})(v_{1}-v_{-})<0\text{,}$ (131)
$v_{+}=-\frac{V_{0}}{V_{0}^{2}+\rho^{2}}+\frac{\rho}{V_{0}^{2}+\rho^{2}}\sqrt{\rho^{2}+V_{0}^{2}-1}\text{,}$ (132)
$v_{-}=-\frac{V_{0}}{V_{0}^{2}+\rho^{2}}-\frac{\rho}{V_{0}^{2}+\rho^{2}}\sqrt{\rho^{2}+V_{0}^{2}-1}<0,$ (133)
whence
$v_{1}<v_{+}\text{.}$ (134)
One can check easily that $v_{+}<1$.
The requirement $v_{+}\geq 0$ gives us
$\rho^{2}\geq 1\text{.}$ (135)
In particular, for the Schwarzschild metric $\rho=0$, $v_{+}<0$ and the PP is impossible as it should be.
Eq. (135) can be rewritten in the form
$\left|q_{1}\right|\varphi\geq\frac{X_{0}}{m_{0}}m_{1}\text{.}$ (136)
Thus in the present section we suggested description of decay and PP in kinematic language by analogy with the rotating case wald . However, there is qualitative difference. For the existence of the PP, in the rotating case the relative velocity between a new fragment and particle 0 should be quite high, the particle being ultrarelativistic wald . In our case, there is no restriction on velocity that can be arbitrarily low. Moreover, instead of the lower bound typical of the rotating metric and process with neutral particles, now there exists the upper bound (134).
XI Scenario B and the maximum of efficiency
Both scenarios I and II imply that the parent particle 0 moves from infinity towards a black hole. Meanwhile, of interest is also another situation when particle 0 bounces back from its turning point and only afterwards decays to particles 1 and 2. Particle 1 flies towards a back hole, particle 2 moves in the outward direction. We call this scenario B (the first letter of the word ”bounce”). Then, the formulas for the energies read
$E_{2}=q_{2}\varphi+\gamma_{02}\frac{m_{2}}{m_{0}}(X_{0}+v_{2}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}})\text{,}$ (137)
$E_{1}=q_{1}\varphi+\gamma_{01}\frac{m_{1}}{m_{0}}(X_{0}-v_{1}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}})\text{.}$ (138)
If we assume that in the frame comoving with particle 0, it is particle 2 that moves outwardly but is drifted in the inward direction due to motion of particle 0, then instead of (118) we have
$V_{2}=\frac{V_{0}+v_{2}}{1+v_{2}V_{0}}\text{.}$ (139)
One can check using (137) and (125) that (139) is indeed satisfied.
Here, as before, particle 2 moves to infinity (now in the same direction as particle 0) and particle 1 moves in the inward direction (opposite to particle 0). However, now the signs before the radicals are opposite to those in (127), (128). The condition for the PP now gives us
$\gamma_{01}\frac{m_{1}}{m_{0}}(X_{0}-v_{1}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}})<\left|q_{1}\right|\varphi$ (140)
instead of (129). This entails
$v_{1}^{2}-\frac{2v_{1}V_{0}}{V_{0}^{2}+\rho^{2}}+\frac{1-\rho^{2}}{V_{0}^{2}+\rho^{2}<0}=(v_{1}-v_{-})(v_{1}-v_{+})<0\text{,}$ (141)
$v_{\pm}=\frac{V_{0}}{V_{0}^{2}+\rho^{2}}\pm\frac{\rho}{V_{0}^{2}+\rho^{2}}\sqrt{\rho^{2}+V_{0}^{2}-1}\text{.}$ (142)
It is easy to check that $v_{+}<1$. Here, there are two different cases.
If
$1-V_{0}^{2}\leq\rho^{2}\leq 1,$ (143)
then $v_{-}\geq 0$ and
$v_{-}<v_{1}<v_{+}\text{.}$ (144)
If
$\rho^{2}>1,$ (145)
then $v_{-}<0$ and
$v_{1}<v_{+}\text{,}$ (146)
so we have an upper bound on $v_{1}$.
If $\rho\rightarrow\infty$, $v_{+}\rightarrow 1$, so actually bound (146) is satisfied automatically.
One can also choose TP3 as a scenario intermediate between I and B. Then, again signs can be arranged according to (108), (109) with the same conclusions about restrictions on the velocity of particle 1 that are necessary for the PP to exist.
The significance of scenario B consists in that it enables us to attain the maximum efficiency $\eta=\frac{E_{2}}{E_{0}}$ of the process. Indeed, in this case the energy of particle 2 is given by eq. (137) with sign ”plus” before the radical.
It is seen from (106) or (137) that $E_{2}$ is monotonically decreasing function of $r$ since $\varphi=\frac{Qq_{2}}{r}$ with $q_{2}>0$ and $\frac{dN}{dr}>0$. Therefore, the most possible maximum of $E_{2}$ is attained when decay happens near the horizon. One should bear in mind that in this case $X_{0}$ should have the order $N_{d}$, so this is possible for near-critical particle 0 only.
However, some important reservations about near-horizon decay are in order. For a nonextremal black hole there is a potential barrier of a finite height that prevents a near-critical particle with $X=(N_{d})$ in the near-horizon region from approaching the horizon. Moreover, if a particle is exactly critical, near the turning point $r_{t}$ from which a particle bounces back, $X^{2}=O(r_{t}-r_{+})^{2}$ near the horizon. But for nonextremal black holes $N^{2}\sim(r-r_{+})$ near the horizon, so $X^{2}\sim N^{4}$. As a resut, it is seen from (6) that the condition $P^{2}\geq 0$ cannot be satisfied and a particle cannot penetrate into the near-horizon region at all. If it is not exactly critical but near-critical, a particle is able to move in the near-horizon region but such a particle can exist only between the horizon and the turning point, so it cannot arrive from infinity anyway. The situation is completely similar to that for rotating black holes gp , gp42 , prd , nearoz . However, if a black hole is extremal (or at least near-extremal) this becomes possible - see Sec. III of Ref. rn for more details about motion near turning points in the case of the extremal RN black hole.
XII Relation to the BSW effect
Up to now, we discussed the standard Penrose process based on particle decay. Meanwhile, the aforementioned properties are applicable to the collisional version of the PP as well, when particles 1 and 2 collide to produce particles 3 and 4. This is due to the fact that particles 1 and can be considered as a combined one with characteristics obeying (10) - (13) and $m_{0}$ equal to the energy $E_{c.m.}$ in the center of mass frame tz . The BSW process and properties of debris were considered before in somewhat different approach in which consideration was restricted from the very beginning to the immediate vicinity of the horizon and approximate formulas were used upper . Now, it is instructive to compare it with the present approach when one starts from exact formulas from the very beginning.
We are mainly interested in the situation when collisions lead to the BSW effect. To make comparison possible, we assume that the RN black hole is extremal and consider pure radial motion like in rn . For the extremal RN black hole, the Coulomb potential $\varphi=1-N$. Let particle 1 be critical. By definition, this means that on the horizon $X=0$. Then, in this setting, $X_{1}=E_{1}N$. We want to show that our exact formulas in the limit when $N_{c}\rightarrow 0$ turn into the results of rn .  (Here subscript ”c” denotes the point of collision.) It is sufficient to trace correspondence with eq. (26) of rn which is the key point of the analysis there.
According to the BSW effect ban and its electric counterpart jl , the energy $E_{c.m.}$ in the center of mass of two colliding particles grows unbounded if (i) one of particles (say, 1) is critical, (ii) collision happens near the horizon, when $N\rightarrow 0$. In our context,
$m_{0}^{2}\approx\frac{\beta}{N}\text{,}$ (147)
where $\beta$ is a constant centr equal to
$\beta=2\left(X_{2}\right)_{c}A\text{, }A=E_{1}-\sqrt{E_{1}^{2}-m_{1}^{2}}\text{.}$ (148)
This formula can be obtained in the near-horizon limit directly from (119), if it is applied to particles 1 and 2. As a result,
$\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}\approx X_{0}-\frac{\beta N}{2X_{0}}\text{,}$ (149)
where $X_{0}=X_{2}+E_{1}N\approx X_{2}$ near the horizon.
Particle 4 falls in a black hole, particle 3 is near-critical particle. Similarly to rn , we define
$q_{3}=E_{3}(1+\delta)\text{,}$ (150)
where
$\delta=CN_{c}\ll 1\text{,}$ (151)
subscript ”c” means the point of collision, $C$ is a constant.
Then, $X_{3}$ and $X_{4}$ are given by our formulas in which 1 should be replaced by 4 and 2 should be replaced by 3.  We have
$X_{3}=E_{3}-q_{3}+q_{3}N=E_{3}(N-CN_{c})\text{,}$ (152)
$X_{3}(N_{c})=E_{3}N_{c}(1-C)\text{,}$ (153)
$P_{3}(N_{c})=N_{c}\sqrt{E_{3}^{2}(1-C)^{2}-\beta}\text{.}$ (154)
From the other hand, it follows from (122) with $\gamma_{01}$ replaced by $\gamma\equiv\gamma_{12}$ that
$\gamma m_{1}m_{2}\approx\frac{\left(X_{2}\right)_{c}(E_{1}-\sqrt{E_{1}^{2}-m_{1}^{2}})}{N_{c}}\text{,}$ (155)
$b_{3}=m_{0}^{2}+m_{3}^{2}-m_{4}^{2}\text{,}$ (156)
$d=b_{3}^{2}-4m_{0}^{2}m_{3}^{2}.$ (157)
Taking into account (49) for $N_{c}\ll 1$, we obtain
$E_{3}(1-C)=\frac{m_{3}^{2}}{2A}+\frac{1}{2}A\text{.}$ (158)
This agrees completely with eq. (26) rn , so further analysis from the aforementioned paper applies. See also upper . If particle 1 moves from the vicinity of the horizon in the outward direction (so-called Schnittman process revised ), consideration runs along the same lines.
XIII Summary and Conclusions
Thus we considered the Penrose process for motion of all three particles within the same plane. We relied on exact formulas for characteristics of daughter particles in terms of a parent one. In doing so, we gave full classification of possible scenarios and derive the bounds on possible values of angular momenta of daughter particles. We selected those scenarios for which ejection occurs along the trajectory of a parent particle. Significance of scenarios of such a type consists in that we obtain maximum (minimum) value of energy. This is confirmed by direct computation of extrema of the energy of corresponding particles with respect to the angular momentum of one of them.
We formulated the results in two forms. The first one includes dependence of the outcome of masses of particles. The second form expresses them in terms of velocities and gamma factors of relative motion between a parent particle and daughter ones. The corresponding formulas obtained for a static charged metric are similar to whose obtained in wald for rotating metrics and neutral particles. Meanwhile, there are some essential differences. For the Penrose process to occur in the Kerr or other rotating metric, there is a lower bound on velocity of a daughter particle in the center of mass frame, this bound being rather high wald . This creates obstacles for using this process. Meanwhile, in our case, the situation is more diverse and there are scenarios in which the velocity obeys the upper (not lower) bound, so the Penrose process can be realized more easily.
Although we discussed decay, the results are applicable to other reactions between particles including the collisional Penrose process in which particles 1 and 2 collide to create new particles 3 and 4. In particular, this can lead to the BSW effect. We traced how the current approach based on exact formulas agrees with the previous one used before for description of the aforementioned effect.
In our consideration, we implied that the metric is the Reissner-Nordström one. However, the approach can be applied to a more general metrics of type (1).
Our goal consisted in the derivation and analysis of possible types of scenarios in the simplest case of processes with charged particles thus developing general formalism. This implied consideration of a static metric and electric field. The next step is supposed to be inclusion of rotating, electric charge and a magnetic field altogether into consideration. We hope that the corresponding formalism will be useful for solving realistic problems in astrophysics. However, this needs separate treatment.